# Shallow learning enables real-time inference of molecular composition from spectroscopy of brain tissue

### v3

@@ Abstract
Optical imaging modalities such as near-infrared spectroscopy (NIRS) and hyperspectral imaging (HSI) represent a promising alternative for low-cost, non-invasive, and fast monitoring of functional and structural properties of living tissue. Particularly, the possibility of extracting the molecular composition of the tissue from the optical spectra in real-time deems the spectroscopy techniques as a unique diagnostic tool. However, due to a lack of paired optical and molecular profiling studies, building a mapping between a spectral signature and a corresponding set of molecular concentrations is still an unsolved problem. Furthermore, there are currently no established methods to streamline inference of the biochemical composition from the optical spectrum for real-time applications such as surgical monitoring.
In this paper, we develop a technique for fast inference of changes in the molecular composition of brain tissue. We base our method on the Beer-Lambert law to analytically connect the spectra with concentrations and use a deep-learning approach to significantly speed up the concentration inference compared to traditional optimization methods. We test our approach on real data obtained from the broadband NIRS study of piglets’ brains and the HSI dataset of glioma patients. The results demonstrate that the proposed method enables real-time molecular composition inference while maintaining the accuracy of traditional linear and non-linear optimization solvers.
keywords:
optics, photonics, light, hyperspectral, machine learning, neural networks, brain imaging
@@
normal-$\dagger$ The authors contributed equally
@@ 1 Introduction
Various biomedical applications such as laboratory tissue examination or neurosurgery require access to rapid monitoring of intrinsic tissue properties. In particular, neuronavigation would benefit by having structural and functional information about brain matter in real-time [1, 2, 3]. Spatially-resolved maps of the tissue characteristics would allow bypassing invasive disease diagnostics, e.g., biopsy, which halt the operation. Instead, a surgical decision could be made during the operation, reducing its time and preserving a healthy brain.
@@
Optical sensing, such as near-infrared spectroscopy (NIRS) and hyperspectral imaging (HSI), emerge as promising alternatives to address these clinical needs [4, 5]. These techniques can probe biological matter utilizing non-ionizing electromagnetic radiation within the visible and near-infrared ranges. The spectral instrumentation is inexpensive, portable, and allows for continuous monitoring at the bedside [6].
@@
The overarching principle behind spectroscopy-based molecular characterization is to relate the reflection spectrum obtained upon illumination of the tissue surface with its optical properties. The molecules constituting the tissue have unique absorption dependency on the radiation frequency, and thus, the reflection should exhibit molecular absorption signatures in its frequency dependency.
@@
However, several other physical factors contribute to shaping the spectra. These include light scattering on the surface and within tissue volume, autofluorescence, and background illumination [7]. Disentangling these phenomena from a reflectance spectrum is often an ambiguous, ill-posed problem, yet it is crucial for deducing the relation between the reflection and molecular composition. Another complication is a scarcity of available studies in which optical monitoring and biochemical composition analysis are performed simultaneously.
@@
Analytical and statistical approaches exist aiming to unmix optical spectra into the physical phenomena defining the spectra profile under a limited data regime [8, 9, 10, 11, 12, 13]. The first family of methods mitigates the data scarcity by introducing a physical prior to establish the spectrochemical link. Typically, the Beer-Lambert law [14, 5] is used to provide such a link by describing the incoming light’s energy dissipation as an exponentially decaying function of light penetration depth into the tissue:
$\log(I_{R}(\lambda)/I_{0}(\lambda))=-\left[\sum_{i}{c_{i}\mu^{i}_{a}(\lambda)}+s\mu_{s}(\lambda)\right]\cdot l+U$ (1)
@@
Here, $I_{R}(\lambda)$ and $I_{0}(\lambda)$ are the intensities of the reflected and the incoming light, $\mu_{a}$ and $\mu_{s}$ is the absorption and scattering coefficients, the index $i$ denotes the molecule constituting the tissue such as water, fat, hemoglobin or cytochromes, $c_{i}$ denotes the corresponding concentration, and $s$ is the weight of scattering in the total light energy dissipation. The remaining quantities are $\lambda$, which is the light wavelength, $l$ is the light penetration depth, and $U$ describes other physical factors contributing to the energy dissipation of the incoming light or other sources of the optical signal captured by the cameras. Now, in the case of changes in molecular composition over the course of optical monitoring, one can assume that the effects contributing to $U$ either stay constant (e.g., which is a fair assumption for ambient illumination) or change notably less than the total absorption  [^1] In certain scenarios, even changes in scattering are not expected since it is a rather bulk effect dependent on the density of the probed matter rather than molecule-specific one.. Under this assumption, subtraction of two reflection spectra, $\log I_{R}^{2}-\log I_{R}^{1}$, would cancel out or make negligible the term $\delta U=U^{2}-U^{1}$ in the following equation:
$\log(I_{R}^{2}(\lambda)/I_{R}^{1}(\lambda))=-\left[\sum_{i}{\delta c_{i}\mu^{i}_{a}(\lambda)}+\delta(s\mu_{s}(\lambda))\right]\cdot l+\delta U$ (2)
@@
In such a differential form, the Beer-Lambert law can now be used to identify molecular composition. For this, standard least-squares optimization algorithms can be employed to minimize the difference between the real spectra and the spectra obtained from the Beer-Lambert law. As a result of the minimization, the optimal values of the set of concentration changes $\{\delta c_{i}=c_{i}^{2}-c_{i}^{1}\}$ are obtained (alongside the scattering parameters).
Another family of unmixing approaches is based on statistical techniques such as principal component analysis (PCA) [12, 13, 11]. Here, by analyzing a large dataset of reflection spectra using the PCA, one identifies the importance of each wavelength to explain the variance in the dataset. Assuming that from the physics point of view, the variance in the reflection spectra is mainly due to varying molecular absorption, one can try to correlate the output of the PCA analysis with the absorption spectra. In [11], the authors discuss how such correlation can reveal the relative difference in concentration of chromophores.
@@
The overarching drawback of these approaches is the computational time it takes to infer the biochemical composition. For example, the optimization methods take a subsecond time to infer the composition of a single spectrum containing typical for NIRS and HSI number of wavelengths (a few hundred). However, when using modalities providing high spatial resolution imaging for real-time applications, one needs to solve the optimization task in a subsecond time for as many spectra as there are spatial pixels, as every pixel contains its own spectrum. For modalities like HSI, the number of pixels can be easily in the order of $10^{5}$. Providing subsecond timings for simultaneous inference on such an amount of spectra poses a challenge for traditional methods  [^2] Note that under the assumption of negligible contribution from scattering in the modified Beer-Lambert law, the Eq. 2 results in a linear system. Such a system can be efficiently solved using the pseudoinverse, achieving close to real-time computation. However, in general, when including non-linear terms in the Beer-Lambert law formulation (scattering, non-linear absorption effects, etc.), the system of equations can not be solved anymore using pseudoinverse..
The present paper aims to alleviate the computational time limitation of existing methods by proposing a data-driven solution that significantly speeds up the biochemical composition inference. We test two types of deep-learning-based approaches which differ in the way the training dataset is collected: a) directly by generating synthetic spectrum-concentrations pairs from the Beer-lambert law or b) by inferring the molecular composition from real spectra by means of the traditional optimization and then use the pairs of real spectra and inferred concentrations for a neural network training. We evaluate the method on broadband NIRS data (transmission mode) [15] and hyperspectral data (reflection mode) [16] of brain tissue. We achieve a significant speed-up compared to the convex least-squares optimization while maintaining accuracy. To our knowledge, this is the first work that applies a neural-network-based approach to provide real-time inference of cerebral chromophore composition from spectroscopy measurements [17, 18, 19, 20, 21, 22, 23, 24, 25].
Figure 1: The general pipeline describing our approach for inferring molecular concentrations’ changes $\{\delta c_{i}\}$ from changes in the reflection spectra.
@@ 2 Method
Datasets creation.
In our method, we use a supervised data-driven approach by creating a dataset of attenuation-concentration pairs to train a neural network (by attenuation, we imply the logarithm of the reflection: $\Delta A=\log I_{R}^{2}-\log I_{R}^{1}$). We employ two different strategies to create the dataset:
a) The first strategy directly utilizes the Beer-Lambert law to generate the training dataset pairs $(\Delta A,\{\delta c_{i}\})$ with $\Delta A$ being the difference in attenuation between two spectra and $\{\delta c_{i}\}$ the corresponding differences in concentration of chromophores. For each chromophore, we randomly sample values for changes in the molecular composition $\{\delta c_{i}\}$ using the uniform distribution within physiologically plausible ranges.
@@
and therefore also uniformly sample parameters $s_{1},s_{2},b_{1},b_{2}$ within plausible ranges [26]. Subsequently, we input the obtained values into the Beer-Lambert law to obtain synthetic differential attenuations $\Delta A(\lambda)$. This difference in attenuation $\Delta A$ as input and the corresponding $\{\delta c_{i}\}$ as output are then used for training.
@@
b) Given that the distribution of the synthetic spectra obtained according to the strategy described above can be notably different from the distribution of real spectra, this can result in an unsatisfactory network prediction accuracy. Therefore, in addition to (a), we evaluate another strategy for creating a dataset trying to bridge the gap between the physical model and real data. For this, we use traditional least squares minimization to fit the changes in the real reflectance spectrum with the modified Beer-Lambert law. The concentrations $\{\delta c_{i}\}$ found upon the optimization and the corresponding $\Delta A$ constitute the training samples.
Network and optimization details.
The training was performed using a multi-layer perceptron (MLP) neural network for both approaches. The network takes as input a one-dimensional vector of attenuation difference and outputs molecular concentration changes.
We trained both networks with early stopping when they reached convergence. To find the optimal network architecture, we used the Ray Tune library [27] to validate different MLP architectures (width, number of hidden layers, and activation functions), learning rates, and batch sizes.

### v6

# Learnable real-time inference of molecular composition from diffuse spectroscopy of brain tissue

@@ Abstract
Significance: Diffuse optical modalities such as broadband near-infrared spectroscopy (bNIRS) and hyperspectral imaging (HSI) represent a promising alternative for low-cost, non-invasive, and fast monitoring of living tissue. Particularly, the possibility of extracting the molecular composition of the tissue from the optical spectra deems the spectroscopy techniques as a unique diagnostic tool.
Aim: No established method exists to streamline the inference of the biochemical composition from the optical spectrum for real-time applications such as surgical monitoring. Here, we analyse a machine learning technique for inference of changes in the molecular composition of brain tissue.
Approach: We propose modifications to the existing learnable methodology based on the Beer-Lambert law. We evaluate the method’s applicability to linear and non-linear formulations of this physical law. The approach is tested on data obtained from the bNIRS- and HSI-based monitoring of brain tissue.
Results: The results demonstrate that the proposed method enables real-time molecular composition inference while maintaining the accuracy of traditional methods. Preliminary findings show that Beer-Lambert law-based spectral unmixing allows to contrast brain anatomy semantics such as the vessel tree and tumor area.
Conclusion: We present a data-driven technique for inferring molecular composition change from diffuse spectroscopy of brain tissue, potentially enabling intra-operative monitoring.
@@
$\dagger$ The authors contributed equally
@@ 1 Introduction
Various biomedical applications such as histopathology or neurosurgery require access to rapid monitoring of intrinsic tissue properties. In particular, neuronavigation would benefit by having structural and functional information of brain tissue in real-time [1, 2, 3]. Spatially-resolved maps of the tissue characteristics would allow bypassing invasive disease diagnostics, e.g., biopsy, which halt the operation. Instead, a surgical decision could be made during the operation, reducing its time and preserving a healthy brain.
@@
Diffuse optical modalities, such as near-infrared spectroscopy and hyperspectral imaging, emerge as promising technologies to address these clinical needs [4, 5]. These techniques can probe biological matter utilizing non-ionizing electromagnetic radiation within the visible and near-infrared ranges. The spectral instrumentation is inexpensive (compared to other modalities such as MRI), allows for continuous tissue monitoring, and can be easily adapted for the complex context of the operating room [6].
@@
The overarching principle behind spectroscopy-based molecular characterization is to relate the reflection spectrum obtained upon illumination of the tissue surface with its optical properties. The molecules constituting the tissue have unique absorption dependency on the optical radiation frequency, and thus, the reflection should exhibit molecular absorption signatures in its frequency dependency.
@@
However, several other physical factors contribute to shaping the measured reflectance spectra. These can include light scattering on the surface and within tissue volume, autofluorescence, tissue inhomogeneity, and background illumination [7]. Disentangling these phenomena from a reflectance spectrum is often an ambiguous, ill-posed problem, yet it is crucial for deducing the relation between the reflection and molecular composition. Another complication is a scarcity of available studies in which optical monitoring and quantitative biochemical composition analysis are performed simultaneously.
@@
Analytical and statistical approaches exist aiming to unmix optical spectra into the physical phenomena defining the spectra profile under a limited data regime [8, 9, 10, 11, 12, 13, 14]. A large number of methods mitigate the data scarcity by introducing a physical prior to establish the spectrochemical link. Typically, the modified Beer-Lambert law [15, 16] is used to provide such a link by describing the incoming light’s energy dissipation as an exponentially decaying function:
$\log(I_{R}(\lambda)/I_{0}(\lambda))=-\left[\sum_{i}{c_{i}\mu^{i}_{a}(\lambda)}+s\mu_{s}(\lambda)\right]l+U$ (1)
@@
Here, $I_{R}(\lambda)$ and $I_{0}(\lambda)$ are the intensities of the reflected and the incoming light, $\mu_{a}$ and $\mu_{s}$ are the absorption and scattering coefficients, the index $i$ denotes the molecule constituting the tissue such as water, fat, hemoglobin and cytochromes, $c_{i}$ denotes the corresponding concentration (e.g. as volume fraction), and $s$ is the weight of scattering in the total light energy dissipation. The remaining quantities are $\lambda$, which is the light wavelength, $l$ is the light pathlength [^1] Even though several works demonstrate the importance of wavelength-dependent definition of the pathlength [17], in what follows, for simplicity, we opted for constant pathlength independent of the wavelength., and $U$ describes other physical factors contributing to the energy dissipation of the incoming light or other sources of the optical signal captured by the light detectors or cameras.
Typical molecules whose changes in concentration are inferred include oxyhemoglobin, deoxyhemoglobin, and cytochrome-c-oxidase (CCO)[5]. Measurement of the former two chromophores can reveal e.g. oxygenation status of the brain, and can help determine hypoxic or hyperoxic conditions. CCO is a fundamental metabolic molecule correlated to ATP production during cellular respiration, which has previously complemented obtained hemodynamic information in various applications[5, 18].
Now, in the case of changes in molecular composition over the course of optical monitoring, one can assume that the effects contributing to $U$ either stay constant (e.g., which is a fair assumption for ambient illumination) or change notably less than the total absorption [^4] In certain scenarios, even changes in scattering are not expected since it is a rather bulk effect dependent on the density of the probed matter rather than molecule-specific one.. Under this assumption, subtraction of two reflection spectra, $\log I_{R}^{2}-\log I_{R}^{1}$ (traditionally, these have been two different points in time, but this could also be true for different points in space), would cancel out or make negligible the term $\delta U=U^{2}-U^{1}$ in the following equation:
$\log(I_{R}^{2}(\lambda)/I_{R}^{1}(\lambda))=-\left[\sum_{i}{\delta c_{i}\mu^{i}_{a}(\lambda)}+\delta(s\mu_{s}(\lambda))\right]l+\delta U$ (2)
@@
In such a differential form, the modified Beer-Lambert law can now be used to identify molecular composition. For this, standard least-squares optimization algorithms (or non-negative matrix factorization [10, 19]) can be employed to minimize the difference between the real spectra and the spectra obtained from the modified Beer-Lambert law. As a result of the minimization, the optimal values of the set of concentration changes $\{\delta c_{i}=c_{i}^{2}-c_{i}^{1}\}$ are obtained (alongside the scattering parameters).
@@
The overarching drawback of this approach is the computational time it takes to infer the biochemical composition. For example, the optimization methods take a subsecond time to infer the composition of a single spectrum containing a number of wavelengths typical for bNIRS and HSI (a few hundred). However, for real-time applications particularly in the case of HSI modality, one needs to solve the optimization task in a subsecond time for as many spectra as there are spatial pixels, as every pixel contains its own spectrum. The number of pixels on a hyperspectral image can be easily in the order of $10^{5}$-$10^{6}$. Providing subsecond timings for simultaneous inference on such an amount of spectra poses a challenge for traditional methods.
Contributions.
There are numerous studies analyzing the application of machine learning methods to achieve fast inversion of the physical models based on Beer-Lambert law or Monte-Carlo simulations [20, 21, 22, 23, 24, 25, 26]. Predominantly they imply training a machine learning model on synthetic data generated by following the chosen physical formalism, and then evaluating the trained model on real spectra. While proven to work for the use cases mentioned in the cited works, this approach might be inferior as synthetic data generators likely underestimate the complexity of real data. To mitigate this, we tested different strategies for model training using only the synthetic data as in previous works or incorporating real data via traditional optimization in the training procedure.
Second, we test the proposed method on its ability to approximate physical models of varying complexity: linear (absorption only) and non-linear (absorption combined with scattering). While evaluated independently in previous works, here we also analyse our approach to explicitly compare both models in terms of spectral fit. This comparison is motivated by a desire to elucidate the conditions under which the linear model (that can be easily solved using, e.g. pseudoinverse) is appropriate for describing the light-brain matter interaction process, and where inclusion of the scattering is necessary.
Third, given that the work is carried out within the HyperProbe project [27] aiming to achieve real-time brain tissue monitoring, the present paper evaluates the computational timing for the biochemical composition inference across different methods and hardware platforms. To reinforce the comparison, in contrast to previous works manually choosing the hyperparameters’ values of the machine learning methods, we used the AutoML technique [28] to identify the most optimal hyperparameters set.
Finally, to our knowledge, this is the first work that applies a neural-network-based approach to provide real-time inference of chromophore composition from in-vivo brain tissue spectroscopy measurements [29, 30, 31, 32, 33, 34, 35, 36, 37]. We evaluate and discuss applicability of the method on broadband NIRS (transmission mode) [38] and hyperspectral (reflection mode) [39] measurements of brain tissue.
@@ 2 Method
Figure 1: The general pipeline describing the learnable approach for inferring concentrations’ changes of molecules such as reduced and oxidized cytochrome-c-oxidase, oxy- and deoxy-hemoglobin $\{\delta c_{i}\}=\{\delta redCCO,\delta oxCCO,\delta HHb,\delta HbO_{2},etc\}$. The pipeline involves training on a dataset which is generated by the means of a modified Beer-Lambert model. According to the model, the light reflection $I_{R}$ is shaped by the absorption $\mu_{a}$ and scattering $\mu_{s}$ phenomena.
2.1 Datasets creation.
In our method, Fig. 1, we use a supervised data-driven approach by creating a dataset of attenuation-concentration pairs to train a neural network (by attenuation, or more precisely the change of it, we imply the logarithm of the reflection: $\Delta A=\log I_{R}^{2}-\log I_{R}^{1}$). We employ two different strategies to create the dataset:
a) The first strategy directly utilizes the modified Beer-Lambert law to generate the training dataset pairs $(\Delta A,\{\delta c_{i}\})$ with $\Delta A$ being the difference in attenuation between two spectra and $\{\delta c_{i}\}$ the corresponding differences in concentration of chromophores, Fig. 2 (a). For each chromophore, we randomly sample values for changes in the molecular composition $\{\delta c_{i}\}$ using the uniform distribution within physiologically plausible ranges. These ranges were determined based on values typically used in the literature, and further details are provided in the appendix.
@@
and therefore also uniformly sample parameters $s_{1},s_{2},b_{1},b_{2}$ within plausible ranges [40]. Subsequently, we input the obtained values into the modified Beer-Lambert law to obtain synthetic differential attenuations $\Delta A(\lambda)$. This difference in attenuation $\Delta A$ as input and the corresponding $\{\delta c_{i}\}$ as output are then used for training.
@@
b) Given that the distribution of the synthetic spectra obtained according to the strategy described above can be notably different from the distribution of real spectra, this can result in an unsatisfactory network prediction accuracy. Therefore, in addition to (a), we evaluate another strategy for creating a dataset trying to bridge the gap between the physical model and real data, Fig. 2 (b). For this, we use traditional least squares minimization to fit the changes in the real reflectance spectrum with the modified Beer-Lambert law. The concentrations $\{\delta c_{i}\}$ found upon the optimization and the corresponding $\Delta A$ constitute the training samples.
2.2 Network and optimization details.
The training was performed using a multi-layer perceptron (MLP [41]) neural network for both approaches. The network takes as input a one-dimensional vector of attenuation difference and outputs molecular concentration changes.
We trained both networks with early stopping when they reached convergence. To find the optimal network architecture, we used the Ray Tune library [28] to validate different MLP architectures (width, number of hidden layers, and activation functions), learning rates, and batch sizes. More details regarding the networks and the training procedure can be found in the appendix.
@@
2.3 Data.
For our experiments, we applied two types of Beer-Lambert law formulation, with and without scattering, to two types of spectral datasets: broadband NIRS data for which the spectra were measured in light transmission mode [38] and hyperspectral data which were obtained in non-contact reflection mode [39].
Figure 3: Optical spectra from the broadband bNIRS study in [38] before, during and after inducing HI in the piglet’s brain (left). On the right, predictions of the molecular concentration change $\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$ over the course of the optical monitoring (for the Beer-Lamber model without scattering). The vertical lines denote the time points corresponding to the normalized reflection spectra on the left.
2.4 Broadband NIRS.
The first dataset is composed of broadband NIRS spectra from a study analyzing 27 piglets’ brains in which a hypoxia-ischemia (HI) state was induced [38]. The piglets were monitored for several hours, during which the carotid arteries were surgically isolated, and a stepwise hypoxia took place for 15-20min. This produced a significant hypoxic-ischaemic effect that changed the metabolic status of the brain and, in some instances, caused further brain injury. The details of the intervention protocol are described in [38]. The optical device used in the study utilizes a miniature light source and a customized high-throughput miniature spectrometer, connected to high numerical aperture optical fibers. The measurements contain around eight thousand spectra per piglet. The distance between each measured time point is between 10.0 and 10.5 seconds. We use the first thousand measurements, i.e. we only consider the first $\sim$2.5 hours of measurement. For all piglets, this is sufficient to observe HI and recovery after HI. As Eq. 2 requires defining a baseline spectrum, analogously to [38], we used a spectrum at the very beginning of optical monitoring (i.e., before HI) for the baseline. We normalized the spectra with respect to dark noise and white reference. The normalized bNIRS spectra before and after the intervention inducing hypoxia are shown in Fig. 3 (left), and predictions of the concentrations change over the course of the optical monitoring are shown in Fig. 3 (right). Out of the 27 piglets in the dataset, 25 had data available during HI, such that 19 were used for training, two for validation, and four for testing. For this dataset, we predict three types of molecules: oxyhemoglobin, deoxyhemoglobin, and differential cytochrome-c-oxidase (CCO) [^2] As the total CCO concentration may be assumed to not change within a few hours, the oxidized-reduced difference spectrum $\mu_{a}^{\text{diffCCO}}$ may be used to infer changes of both oxidized and reduced CCO [18]. , i.e. $\{\delta c_{i}\}=\{\delta c_{HbO_{2}},\delta c_{HHb},\delta c_{\text{diffCCO}}\}$, where $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$. We neglected the potential contribution to the spectra from water and fat due to their minimal change in concentrations during the 2.5 hours of monitoring [38, 43]. Note that we assume unitary pathlength in our experiments, which results in units of [mM/cm] and [1/cm] for the inferred concentrations. In the NIR range, it has been shown that the pathlength is semi-constant [38], which effectively leads to a simple rescaling in our concentrations when using this assumption. This can also be observed when comparing the inferred exemplary concentrations from Fig. 3 (right) with pathlength-corrected concentrations in previous work [38].
2.5 Hyperspectral data.
The second dataset we used consists of hyperspectral data from the Helicoid project [39]. The Helicoid dataset comprises brain HSI images obtained in surgical conditions from 22 patients diagnosed with glioma. The optical instrumentation is based on pushbroom technique and a silicon CCD detector array as a camera. The HSI images provide a high spectral resolution of 826 bands spread between 400 and 1000 nm and a 2D spatial resolution of a few hundred pixels in each dimension. The images were also expert-annotated into three tissue classes: normal and tumor tissues, as well as blood vessels. Typical hyperspectral image and corresponding spectra are shown in Fig. 4. Different from the bNIRS dataset, we used a spectrum of the pixel belonging to the blood vessel class as a baseline spectrum [^4] The blood vessel was used as a reference for a couple of reasons. The blood vessel is clearly distinguishable from the other two tissue types, tumor and non-tumor tissue. These two types of tissue are highly heterogeneous, e.g. within a pixel area, they can have small capillaries, leakage of blood, agglomeration of dead cells, etc. In contrast, the blood vessel pixels, especially the ones belonging to large arteries, are less heterogeneous. Moreover, it is assumed that blood vessels do not possess cytochrome molecules. Thus, it is a better reference when one sets a goal of detecting the presence of cytochromes in the brain matter. . We then subtracted the baseline spectrum from all other spectra in the same image. In other words, we performed the differential spectroscopy not in time but in space.
Figure 4: Optical spectra from the HSI study of patients diagnosed with glioma [39] for different tissue types: tumor, normal tissue, and blood vessels (left). On the right is a typical RGB image of the brain surface, which is obtained from the HSI volume. The dots correspond to the spectra on the left image. The black circles on the RGB image are rubber rings that surgeons used to mark tumomr and healthy tissues.
Figure 5: Examples of HSI images of two patients, shown with respective patient ID, from the Helicoid dataset [39] (left). Each pixel in the shown 2D image possesses a spectral signature with 826 bands. From this signature, we predict the molecular concentration change for hemodynamic $\delta c_{HbT}=\delta c_{HbO_{2}}+\delta c_{HHb}$ (middle), and metabolic $\delta c_{\text{diffCCO}}=\delta c_{oxCCO}-\delta c_{redCCO}$ characterization (right). Here, we use the Beer-Lambert model with scattering, as it provides a closer fit to real spectra than the model without scattering. We observe that performing the spectral unmixing on the HSI measurement of brain tissue allows us to better contrast the vessel tree (middle) and tumor area (right) than on the RGB image.
Besides predicting oxyhemoglobin $\delta c_{HbO_{2}}$ and deoxyhemoglobin $\delta c_{HHb}$, we again infer the differential cytochrome-c-oxidase concentration due to its role in capturing oxidative metabolic activity. We separately predicted oxidized cytochrome-c-oxidase and reduced cytochrome-c-oxidase, as the total CCO concentration may not be assumed to remain constant in space. We also predict water and fat since, for these molecules, one cannot assume minimal concentration change across different tissue types as in the case of the bNIRS spectra. For reference, the absorption spectra can be found in the appendix.
Fig. 5 showcases examples of molecular inference for the HSI images from the Helicoid dataset. Out of the nine patients with glioblastoma in the dataset, six with distinct class labeling were chosen, and three patients were used for training, one for validation, and two for testing. Note that patients might have multiple images taken, and different images from the same patient were assigned to the same training, validation, and test set to avoid set contamination. Therefore, the training set consists of five, the validation set of one, and the test set of three images.